Q=400-2P+4i and Q=100+30

1
Expert's answer
2021-11-19T11:02:27-0500

At Equilibrium; Quantity supplied = Quantity Demanded

"400-2P+4i=100+3P"

Assuming i=25

Then

"400-2P+4*25=100+3P"

"400-2P+100=100+3P"

"500-100=3P+2P"

"400=5P"

"P=80"

"Q=400-2(80)+100"

"Q=340"